A stock with a beta of 1.2 just paid a dividend of $0.75 that is expected to grow at 7%. If the risk-free rate is 3% and the market risk premium is 5.5%, what should be the price of the stock in five years? A. $28.85 B. $43.29 C. $30.87 D. $40.46
Answer #1

Required return=risk free rate+Beta*market risk premium

=3+(1.2*5.5)=9.6%

Current price=D1/(Required return-Growth rate)

=(0.75*1.07)/(0.096-0.07)

=30.8653846

P5=Current price*(1+Growth rate)^5

=30.8653846*(1.07)^5

=$43.29(Approx)